A sample of oxygen gas occupies a volume of 7.93 L at 59 °C and 843 mm Hg. The volume of the gas sample increases to 8.96 L while its pressure increases to 1.13 x 10^3 mm Hg. If the ideal gas law is obeyed, the temperature of the gas sample at the new volume and pressure is _____ °C.


1
Expert's answer
2022-03-31T05:12:05-0400

p1V1/T1 = p2V2/T2

T2 = p2V2T1/p1V1 = (8.96 L)×(1.13×103 mm Hg)×(59 + 273.15 K) / (7.93 L)×(843 mm Hg) = 503 K or 230°C